OpenStudy (anonymous):

How do you simplify the square root of 180 and how do you know that it is in simplest form?

Parth (parthkohli):

You factorize it.

OpenStudy (campbell_st):

look for a square number that is a factor of 180 you could have 4 x 45, 9 x 20, or 36 x 5 once you have the factors the question can be written as \[\sqrt{180} = \sqrt{36 \times 5}\] I'll let you finish the answer